OpenStudy (anonymous):

I WILL MEDAL AND FAN THE BEST ANSWERS!!!!!! PLZZZZZ HELP ME :) What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. -5 m/s2 B. +5 m/s2 C. -25 m/s2 D. +25 m/s2

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

You can apply this formula \(a=\Large \frac{v_{final}-v_{initial}}{t}\)

OpenStudy (anonymous):

as @tinybookworm says, \[\frac{ 0-25 }{ 5 }= -5 m/s2\]